According to Masterfoods, the company that manufactures M&M’s, 12% of peanut M&M’s are brown, 15% are yellow, 12% are red, 23% are blue, 23% are orange and 15% are green. You randomly select six peanut M&M’s from an extra-large bag of the candies. (Round all probabilities below to four decimal places; i.e. your answer should look like 0.1234, not 0.1234444 or 12.34%.)

Compute the probability that exactly five of the six M&M’s are red

Compute the probability that four or five of the six M&M’s

Compute the probability that at most five of the six M&M’s are red.

Compute the probability that at least five of the six M&M’s are red.

If you repeatedly select random samples of six peanut M&M’s, on average how many do you expect to be red? (Round your answer to two decimal places.)

red M&M’s With what standard deviation? (Round your answer to two decimal places.)
Answer #1

here this is binomial distribution with parameter n=6 and p=0.12

1)

probability that exactly five of the six M&M’s are red =P(X=5)=6C5(0.12)5(0.88)1 =0.0001

2) probability that four or five of the six M&M’s =P(X=4)+P(X=5)=6C4(0.12)4(0.88)2+6C5(0.12)5(0.88)1=0.0025

3)

probability that at most five of the six M&M’s are red =P(X<=5)=1-P(X=6)=1.0000

4)

probability that at least five of the six M&M’s are red =P(X>=5)=P(X=5)+P(x=6)=0.0001

5) expect to be red =np=6*0.12 =0.72

6) standard deviation =sqrt(np(1-p))=sqrt(6*0.12*0.88)=0.80